OpenStudy (babystar16):

find the equation of a line with slope -3 and goes through the point (-5,2)

OpenStudy (anonymous):

you have to use the slope-point equation of a line. when you're given a point and a slope. do you know this?

OpenStudy (babystar16):

kinda i missed some of this stuff and yah im in geomerty and mathmodels so yah Help plzz

OpenStudy (nubeer):

m=-3 , (-5,2) y-y1 = m(x-x1) y1 = 2 , x1 = -5 just plug in values .. you will get the equation

OpenStudy (anonymous):

ok so you've been answered already. hope you understand it :)